OpenStudy (anonymous):

b) If you have a rectangular rug whose perimeter is 26 feet and the length is three feet more than the width, what are the dimensions of the rug? (P=2w+2l)

OpenStudy (anonymous):

length=8 width=5

OpenStudy (anonymous):

26=2(5)+2(8) 26=10+16 26=26 IS THIS RIGHT?